Grade 12th passPhysical Chemistry

What is the value of velocity of an electron in 2nd shell of li2+ ion?

Ankita das
8 Years agoGrade 12th pass

1 Answer

Nikhil Goyal
ApprovedApproved Tutor Answer8 Years ago
We will use formula z/n×2•1888×10^6m/sz=3,n=2,3/2×2•1888×10^6=3•2832×10^6m/s✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓✓